Information for the equipment purchase of Eshy25 Company for 20X6 is as follows:

Date of purchase January 1
Purchase of machine $5,000,000
Useful life 5 years

It is Eshy's policy to use double‐declining balance for the first two years of the asset's life and switch to straight‐line depreciation from the third year onward. How much is the accumulated depreciation for the December 31, 20X8 balance sheet?

Group of answer choices

a. 3,800,000

b. 3,600,000

c. 3,000,000
Answer #1

EF total dep B year 1 2 Machine Value 5,000,000.00 3 Depreciation 2,000,000.00 4 Balancing Value 3,000,000.00 year 2 year3 3,

2 Machine Value 3 Depreciation 4 Balancing Value B year 1 5000000 =+B2*40% =+B2-B3 year 2 =+B4 =+C2*40% =+C2-C3 DE year3 tota

please ask if any qury.